OpenStudy (anonymous):

Can someone help me please? tan^2(x) +5 tan(x) +3=0 Approximate to four decimal places the solutions for x, 0 less than or equal to x<2pi

OpenStudy (anonymous):

\[\tan ^{2}x+5\tan x+3=0\] using quadratic equation \[\frac{ -5\pm \sqrt{5^{2}-4(1)(3)} }{ 2(1) }\] so \[\frac{ -5\pm \sqrt{25-12} }{ 2 }\] \[\frac{ -5\pm \sqrt{13} }{ 2 }\] \[\frac{ -5+\sqrt{13} }{ 2 }=-.6972 \] or \[\frac{ -5-\sqrt{13} }{ 2 }=-4.3028\] but it is none of my choices?????

OpenStudy (anonymous):

Those values are for tan x

OpenStudy (anonymous):

So do I need to convert them into sin/cos?

OpenStudy (anonymous):

According to those figures there is no solution in the given interval

OpenStudy (anonymous):

my choices are A) 0.8692, 2.2724, 4.0108, 5.4140 B) 0.9041, 2.2375 C) 1.7992, 2.5327, 4.9407, 5.6743 D) 1.2304, 1.9111, 3.1416